If a 2 + b 2 – c 2 – 2ab = 0 then the point of concurrency of family of straight lines ax + by + c = 0 lies on the line Options y = x y + x + 1 y = - x x + y = 1 Hint

If a2 + b2 – c2 – 2ab = 0 then the point of concurrency of family of straight lines ax + by + c = 0 lies on the line

Options

y = x
y + x + 1
y = - x
x + y = 1

Hint

Grade:

2 Answers

BALAJI ANDALAMALA
8 years ago
Given a^2+b^2-c^2-2ab=0
(a-b)^2-c^2=0
(a-b-c)(a-b+c)=0
a=b+c or b=a+c.

Substitute it in the equation of the line .
(b+c)x+by+c=0
b(x+y)+(x+1)c=0
x+y=0 or x+1=0

y=-x be the required option.

Ketan Agrawal
7 years ago
Your Answer is absolutely right.I am also thinking.But you are great person.So like it.l love you.